OpenStudy (anonymous):

Find the indicated value. Find f(-2) when f(x)=5.5(x+1.5) Is the answer -9.7 ?

OpenStudy (paxpolaris):

\[f(-2)=5.5 \times (-2+1.5)\]

OpenStudy (paxpolaris):

\[=5.5 \times -0.5\]

OpenStudy (paxpolaris):

\[=-\left( {11\over 2}\times {1\over2} \right)=-{11\over4}=-2.75\]

OpenStudy (anonymous):

Thank-you sir